# Inner crust of neutron stars: Unified description of structure and superconductivity

### v1

@@ II.2 Basic equations
As it is explained in Appendix B, within a WS cell the total energy density, Eq. (65), for (1N, 2N, 3N) phases can be written as a sum of the following pieces:
$w_{\mathrm{tot}}=w_{\mathrm{nuc}}+w_{\mathrm{p.surf}}+w_{\mathrm{C+L}}+w_{\mathrm{no}}+w_{\mathrm{e}}+w_{\mathrm{curv}}+w_{\mathrm{sk}}+w_{\mathrm{dc}},$
@@
where $w_{\mathrm{nuc}}$ is the bulk (uniform) energy density of the nuclear cluster. The terms $w_{\mathrm{p.surf}}$ and $w_{\mathrm{curv}}$ are the nuclear surface contributions, where the former describes the planar surface contribution and the latter describes the curvature of the interface. The contribution from the dripped neutron liquid is given by $w_{\mathrm{no}}$ and the kinetic energy of the electron background by $w_{\mathrm{e}}$. The Coulomb energy term $w_{\mathrm{C+L}}$, collects the self-energy of protons in the nuclear cluster and of electrons distributed in the WS cell as well as the proton-electron lattice energy. Finally, the terms $w_{\mathrm{sk}}$ (the contribution from the neutron skin in the planar interface) and $w_{\mathrm{dc}}$ (its correction due to curvature of the interface) provide the energy contribution due to the adsorbed neutrons at the surface of the cluster. Analogously, the energy density for bubble phases is described by Eq. (141).
The basic equations are a set of coupled algebraic equations, which can be obtained from the variational procedure as specified in Appendices B and C. That the variational equations are satisfied implies the system state is in the local energy minimum. In this paper, we will find the solution with $n_{sk}=n_{dc}=0$, which implies neglect the neutron skin energy density correction and also neglect of the width of the thin layer where the neutron number density smoothly transitions between its values inside and outside of the nucleus.
@@
With the mean baryon density $n_{b}$ given as an external parameter, the four unknowns are the proton fraction inside the nucleus, the nucleus (or the bubble) radius, the baryon number density inside the nuclear matter phase and the neutron number density inside the pure neutron liquid:
$\{x,\,r_{N}\,({\rm or}\,r_{B}),\,n,\,n_{no}\}.$ (3)
The full set of basic (variational) equations contains four algebraic equations:
@@
The explicit derivation of Eqs. (4)-(7) is detailed in Appendices B and C. The quantities in Eqs. (4)-(7) are defined for nuclei in Eqs. (88), (95) and (99); (78), (67), (68) and (73); (94), (95) and (96); (106), (107) and (108); for bubbles in Eqs. (162), (169) and (171); (152), (143), (144) and (147); (168), (169) and (170); (178), (179) and (180). Notice that our definitions of the pressure inside the nucleus $P_{i}$ and of the chemical potential of neutrons inside the nucleus $\mu_{ni}$, include both the bulk and the surface contributions, which differs from the notation used by Lim and Holt LimHolt2017 .
In order to solve Eqs. (4)-(7), as a first step we choose the ranges of values of $n$ and $n_{no}$ and represent them as uniform arrays of points. For each pair of values of $n$ and $n_{no}$ from the chosen ranges we find from Eq. (4) the “equilibrium” value of $x=x(n,n_{no})$. In this step we express the volume fraction $u$ from Eqs. (63) and (139) as
@@
In the next step we find the pairs of variables $(n,n_{no})$ that satisfy the rest two of the variational equations, Eqs. (6) and (7). As a result, we obtain two curves on the plane $(n,n_{no})$, which might intersect thereby determining the point of the equilibrium. Having determined the intersection point, we have found all the independent variables of the problem and thus, the solution is accomplished. The calculations including the neutron skin corrections will be presented elsewhere.
III Numerical results
III.1 Comparison with earlier work
III.1.1 Surface tension in LDM
@@
This is in contrast with earlier LDM calculations, where different equations of state were used for describing the bulk matter but with a surface tension given by an analytical expression depending on few parameters ravenhall83 ; lattimer85 ; lorentz93 . For instance, Eq. (9) of LimHolt2017 suggests the following fit
@@
In order to check our numerical algorithm, we compute the inner crust structure using Eq. (9) with the parameters given in Eq. (10) and discarding the curvature contribution. We calculate the transition densities between different structures with numerical error bars and display the results in the third column of Table I (marked as Case 2). A good agreement with the earlier results due to Lim and Holt LimHolt2017 is observed, thereby providing an independent verification of our numerical algorithm. We have also verified that the both parameterizations, Eqs. (11) and (10) with $\sigma_{c}=0$ in both cases, lead to two different solutions of the variational equations with the difference being a small quantity.
As a next step, we include both the planar surface contribution $\sigma_{s}(x)$ and the curvature contribution $\sigma_{c}(x)$, using the numerical data given in Table 3 for $\sigma_{s}(x)$, $\sigma_{c}(x)$ and their derivatives. The transitions between different configurations are listed in the fourth column of Table 1 (marked as Case 3). A comparison between the results shown in the the second and fourth columns of Table 1 (marked as Case 1 and Case 3) suggests that, the inclusion of $\sigma_{c}(x)$ in this paper provides a qualitative difference for the ground state of the inner crust matter, as compared with the ground state predicted earlier in LimHolt2017 .
@@
Table 1: The baryon number densities $\eta\equiv n_{b}/n_{0}$ at which the solution with the minimum energy per baryon changes its symmetry type (between 1N, 2N, 3N, 2B, 3B phases), or $\eta_{\rm uni}^{*}$ at which the uniform nuclear matter is unstable with respect to modulations of the nucleon densities, or $\eta_{\rm p}^{*}$ at which the proton drip occurs, in various cases. Case 1: Output data from Table III of LimHolt2017 . Case 2: Solutions to Eqs. (4)-(7) with $\sigma_{c}=0$ and with $\sigma_{s}$ from Eqs. (9) and (10). Case 3: Solutions to Eqs. (4)-(7) with $\sigma_{s}$ and $\sigma_{c}$ from Table 3. The quantity $\eta_{\rm uni}^{*}$ is found from LimHolt2017 (see Eq. (30) and Table V in LimHolt2017 ). The quantity $\eta_{\rm p}^{*}$ is found from our solution for the workfunction $\Delta\mu_{p}$ shown in Fig. 5.
III.1.2 Towards a realistic structure of the inner crust
The idea that the inner crust might be polycrystalline has been discussed in Refs. KobyakovPethick2015 ; ZhangPethickKobyakov2020 . Lately, Zhang and Pethick ZhangPethick2022 have considered a model of disorder in the pasta phase at a given baryon density (in the absence of the magnetic field), corresponding to a collection of monocrystals of large size with the structure oriented at different angles relative to each other. Still, another option is that the structure orientation varies smoothly in space, similarly as occurs in laboratory liquid crystals ZhangPethickKobyakov2020 . Later, Newton et al. NewtonEtal2022 have suggested that at a constant pressure (in the absence of the magnetic field), several pasta structures are frozen into domains with size of the order of 1-50 times the lattice spacing and over which the local density and electron fraction can vary. The numerical results of Newton et al. NewtonEtal2022 on the basis of a thermodynamic potential $E+PV$ ignoring the magnetic field and entropy, where $E$ is the internal energy per particle, $P$ is pressure and $V$ is specific volume, have suggested that the structure of pasta phases might be amorphous, i.e. without a long-range order.
It seems that a further exploration of thermodynamic properties of the pasta phases on the basis of the complete Gibbs free energy, that includes the entropy and the magnetic induction, is necessary in order to establish definitively, which of the solid phases – amorphous, crystalline or polycrystalline, polymorphous or pure (thermodynamic) – are expected as a result of cooling process in a real neutron star. Notably, the additional complication is that the result of cooling in a real system is affected by many factors including: The order in which the system experiences the liquid-solid phase transition (the formation of cold solid structure from a warm liquid plasma in the inner crust) and the superfluid-superconducting phase transitions, the cooling rate as compared to rate of relaxation of polymorphic crystals to the thermodynamical phase (characterized by the global minimum of the Gibbs free energy), local fluctuations of the proton fraction, structural defects of the solid-state plasma configurations, mechanical vibrations and the local magnetic field.
III.2 Basic solution with corrections for curvature of thin nuclear surface
@@
In this calculation we have incorporated the curvature correction to the surface tension, which is a contribution to the energy due to the principal curvature of a thin nuclear surface without finite-width corrections. Here, in order to use our self-consistent numerical results (as given in Table 3) on the leading (planar surface) contribution $\sigma_{s}(x)$ and the next-to-leading (curvature) correction $\sigma_{c}(x)$, for the surface tension of a thin surface, we represent these functions numerically with the help of spline interpolation – in this way we avoid choosing any fitting parameters such as in Eq. (9). By following the scheme of solution of the variational equations explained above we find the four basic variables, $\{n,x,r_{N}\,{\rm or}\,r_{B},n_{no}\}$ as functions of mean (averaged across the WS cell) baryon density $n_{b}$ scaled by the saturation density in symmetric nuclear matter $n_{0}$ and display them in Figs. 1–4.
@@
Figures 1–4 show, correspondingly, as functions of $n_{b}$: The proton fraction $x$ inside the nucleus (or outside the bubble) in Fig. 1,
@@
the nucleus radius $r_{N}$ (or the scaled bubble radius, $r_{B}\times 10^{-1}$) in Fig. 2,
@@
the local baryon density $n$ inside the nucleus (or outside the bubble) in Fig. 3,
Figure 3: Solution to the basic equations of equilibrium Eqs. (4)-(7): The baryon number density inside the nucleus and outside the bubble as function of the mean baryon density $n_{b}$.
the local number density $n_{no}$ of neutrons outside of the nuclear cluster (or inside the bubble) in Fig. 4.
Figure 4: Solution to the basic equations of equilibrium Eqs. (4)-(7): The baryon number density $n_{no}$ of pure neutron matter outside the nucleus and inside the bubble as function of the mean baryon density $n_{b}$.
III.3 Analysis of the solution
The variables shown in Figs. 1–4 represent a complete and unique solution to the basic equations, however, the appearance of pasta structures is not explicit. As a first step, we compute the difference between the proton chemical potential inside and outside of the nuclear cluster (Eqs. (99), (171) and (100)),
@@ III.3 Analysis of the solution
which represents a potential energy barrier created by neutrons that prevents protons flowing from one nuclear cluster to a neighboring one directly through the pure neutron matter between the clusters.
@@
have been obtained assuming that the second phase is the pure neutron liquid. With this assumption, the surface energy is a function of the proton fraction only inside the nucleus (see Eqs. (67) and (67)). However, in the situation, not considered in this work, when the second phase is a dilute nuclear matter Vinas1998 ; douchin00 ; PethickEtal1995 ; KellerEtal2024 , the present LDM should be extended in order to take into account nonzero proton fractions both inside and outside the nucleus.
@@
Inserting the solutions shown in Figs. 1–4 into the expressions for the energy density, Eqs. (64) and (140), we calculate the total energy density for each of the five structures considered in the crust. Next, we determine the ground-state configuration (with the lowest energy per baryon). Notice that for fixed $n_{b}$, the picture in terms of the energy per baryon is equivalent, up to a constant scaling factor, to the picture in terms of the energy density. As a next step, we find all the possible structures whose energy per baryon differs from the minimum energy per baryon by less than the thermal energy $k_{B}T$.
@@
These results are displayed in Fig. 6, where the minimum energy is shown by black solid line, while the dot markers show the solutions with energy per baryon larger than the minimum energy per baryon by less than the thermal energy $k_{B}T$, for three different temperatures. The vertical blue lines show the critical densities $\eta_{p}^{*}$ and $\eta_{\rm uni}^{*}$. The calculations for $n_{b}/n_{0}>0.6$ have been done on a sparser array of values since the uniform matter is the only possible solution in this case.
@@
Figure 6: (Color online) Thermodynamically allowed pasta configurations at three different temperatures $T$ taking the values $10^{9}$ K (red), $5\times 10^{8}$ K (gold) or $10^{8}$ K (green). Black solid line shows the thermodynamic (the ground-state) configuration characterized by the minimum energy per baryon. The transitions between various pasta phases are seen as jumps of the black solid line, where the almost vertical pieces represent the absence of the data due to the numerical approximations made. The color markers indicate the excited configurations, which satisfy Eqs. (4)-(7), while their energy per baryon is larger than that in the thermodynamic configuration by less than $k_{B}T$. The dash-dotted lines marks the critical value of $n_{b}$ at which uniform nuclear matter becomes unstable with respect to long wavelength hydrodynamic fluctuations (see Eq. (41) for details). The dashed line marks the critical value of $n_{b}$ at which protons start dripping out from the nuclei (see Fig. 5 for details).
Figure 6 suggests that even at $T\sim 10^{8}$ K, the state of a real matter is likely a polymorphic structure, in which locally, nuclear clusters with various types of symmetry (either of $1N,2N,3N,2B,3B$ or uniform) coexist. For instance, for $n_{b}$ below $\sim 0.16n_{0}$, the system energetically favors only 3N phase. However, for $n_{b}$ above $\sim 0.16n_{0}$ and below $\sim 0.297n_{0}$, the system favors both 3N and 2N phases. For $n_{b}$ above $\sim 0.297n_{0}$ and below $\sim 0.345n_{0}$, or above $\sim 0.41n_{0}$ and below $\sim 0.515n_{0}$, the system favors 3N, 2N and 1N phases. Interestingly, for $n_{b}$ above $\sim 0.515n_{0}$, the system favors 3N, 2N, 1N and the uniform phases.
IV Superconductivity in 1-dimensional pasta phase and astrophysical implications
IV.1 Characteristic physical lengths of protons
Let us focus on the 1N phase of the pasta phases. When the 1N phase is superconducting, its ordered state possess the most exotic magnetic properties among other pasta phases due to its extreme anisotropy.
Figure 6 shows that 1N phase is the ground state in a significant portion of the inner crust  [^1] While we have included the curvature contribution to the surface tension, we have ignored the effect of the adsorbed neutrons and other smaller corrections; however, initial calculations by B. K. Sharma (2024, private communication) have shown that changes in the results due to these missed terms are small, in contrast with the case of a weak perturbation of the energy functional due the curvature energy density $w_{\rm curv}$.. As has been pointed out earlier Kobyakov2018 , in 1N phase, the magnetic field effects are especially pronounced when protons are superconducting. In the weak pairing Bardeen-Cooper-Schrieffer approximation, the coherence length, which can be understood as the spatial size of the Cooper pair, reads

### v2

@@ III.2 Basic equations and scheme of numerical solution
As it is explained in Appendix B, within the unit cell the total energy density, Eq. (71), for 1N, 2N and 3N phases can be written as a sum of the following pieces:
$w_{\mathrm{tot}}=w_{\mathrm{nuc}}+w_{\mathrm{p.surf}}+w_{\mathrm{C+L}}+w_{\mathrm{no}}+w_{\mathrm{e}}+w_{\mathrm{curv}},$
where $w_{\mathrm{nuc}}$ is the bulk (uniform) energy density of the nuclear cluster. Analogously, the energy density for bubble phases is described by Eq. (131) and is explained in Appendix C. We neglect a contribution from the neutron skin, a correction due to smearing of the nucleon densities and a correction due to pairing energy.
@@
In the total energy density $w_{\mathrm{tot}}$, the terms $w_{\mathrm{p.surf}}$ and $w_{\mathrm{curv}}$ are the nuclear surface contributions, where the former describes the planar surface contribution and the latter describes the curvature of the nucleus interface. The contribution from the dripped neutron liquid is given by $w_{\mathrm{no}}$ and the kinetic energy of the electron background by $w_{\mathrm{e}}$. The Coulomb energy term $w_{\mathrm{C+L}}$, collects the self-energy of protons in the nuclear cluster and of electrons distributed in the WS cell as well as the proton-electron lattice energy.
@@
The basic equations are a set of coupled algebraic equations, which can be obtained from the variational procedure, as specified in Appendices B-C. Satisfaction of the variational equations implies the system state is in the local energy extremum. With the mean baryon density $n_{b}$ given as an external parameter, the four unknowns are the proton fraction inside the nucleus $x$, the nucleus radius $r_{N}$, the baryon number density inside the nucleus $n$ and the neutron number density inside the pure neutron liquid $n_{no}$:
$\{x,\,r_{N},\,n,\,n_{no}\}.$ (3)
Correspondingly, in the bubble phase the unknowns are the proton fraction outside the bubble $x$, the bubble radius $r_{B}$, the baryon number density outside the bubble $n$ and the neutron number density inside the bubble $n_{no}$.
The full set of basic equations contains four algebraic equations:
@@
The quantities in Eqs. (4)-(7) are defined for the nucleus phases in Eqs. (92), (99) and (103); (82), (73), (74) and (79); (98), (99) and (100); (110), (111) and (112). The corresponding quantities for the bubble phases are defined in Eqs. (150), (157) and (159); (140), (133), (134) and (137); (156), (157) and (158); (166), (167) and (168). Notice that our definitions of the pressure inside the nucleus $P_{i}$ and of the chemical potential of neutrons inside the nucleus $\mu_{ni}$, include both the bulk and the surface contributions, which differs from the notation used by Lim and Holt LimHolt2017 .
In order to solve Eqs. (4)-(7), as a first step we choose the ranges of values of $n$ and $n_{no}$ and represent them as uniform arrays of points. For each pair of values of $n$ and $n_{no}$ from the chosen ranges we find from Eq. (4) the equilibrium value of $x=x(n,n_{no})$. In this step we express the volume fraction $u$ from Eqs. (1) and (129) as
@@
In the next step we find the pairs of variables $(n,n_{no})$ that satisfy the rest two of the variational equations, Eqs. (6) and (7). As a result, we obtain two curves on the plane $(n,n_{no})$, which might intersect thereby determining the point of existence of the numerical solution. Having determined the intersection point, we have found all the independent variables of the problem and thus, the solution is accomplished.
IV Liquid drop model: Numerical results
IV.1 Verification of computational algorithm and choice of surface tension
@@
In order to check our numerical algorithm, we compute the inner crust structure using Eq. (9) with the parameters given in Eq. (10) and discarding the curvature contribution. For convenience, we use throughout the paper the scaled variable
$\eta\equiv\frac{n}{n_{0}}.$ (12)
We calculate the transition densities between different structures with numerical error bars and display the results in the third column of Table I (marked as Case 2). For comparison, we list the earlier results due to Lim and Holt LimHolt2017 in the second column of Table I (marked as Case 1). We observe a good agreement between Case 1 and Case 2 in Table I. This provides an independent verification of our numerical algorithm. We have also verified that two sets of numerical results (with $\sigma_{c}=0$) obtained from two different parameterizations given in Eqs. (11) and (10), are not significantly different.
As a next step, we include both the planar surface contribution $\sigma_{s}(x)$ and the curvature contribution $\sigma_{c}(x)$, using the numerical data given in Table 3 for $\sigma_{s}(x)$, $\sigma_{c}(x)$ and their derivatives. The transitions between different configurations are listed in the fourth column of Table 1 (marked as Case 3). A comparison between the results shown in the the second and fourth columns of Table 1 (marked as Case 1 and Case 3) suggests that, the inclusion of $\sigma_{c}(x)$ in this paper induces a qualitative difference for the ground state of the inner crust matter, as compared with the ground state predicted earlier in LimHolt2017 .
@@
Table 1: The baryon number densities $\eta\equiv n_{b}/n_{0}$ at which the solution with the minimum energy per baryon changes its symmetry type (between 1N, 2N, 3N, 2B, 3B phases), or $\eta_{\rm uni}^{*}$ at which the uniform nuclear matter is unstable with respect to modulations of the nucleon densities, or $\eta_{\rm p}^{*}$ at which the proton drip occurs (see Fig. 5), in various cases. Case 1: Output data from Table III of LimHolt2017 . Case 2: Solutions to Eqs. (4)-(7) with $\sigma_{c}=0$ and with $\sigma_{s}$ from Eqs. (9) and (10). Case 3: Solutions to Eqs. (4)-(7) with $\sigma_{s}$ and $\sigma_{c}$ from Table 3. The quantity $\eta_{\rm uni}^{*}$ is found from LimHolt2017 (see Eq. (30) and Table V in LimHolt2017 ). The quantity $\eta_{\rm p}^{*}$ is found from our solution for the workfunction $\Delta\mu_{p}$ shown in Fig. 5. In this work we focus on Case 3, in which the inclusion of curvature energy leads to dramatic effects, specifically, to disappearance of the bubble phases.
IV.2 Energies of pairing, neutron skin and density smearing
In the main numerical calculations of this paper, we have ignored energies of pairing, neutron skin and density smearing. It appears from calculations by B. K. Sharma Sharma2024 that the neutron skin and density smearing do not significantly affect the pasta phase transition. This is in contrast to the curvature correction, which notably influences the appearance of pasta phases, as can be seen from the results reported in Table I.
We have performed calculations of energies of the pasta phases with the pairing energy density included by adding $w_{p}$ from Eq. (15) (and for the bubble phases, $w_{p}^{\rm bub}$ from Eq. (16), correspondingly) to the total energy density $w_{\rm tot}$ given in Eq. (70). We have found that the pairing has no effect on the pasta phase transition densities within the actual precision of our computation. The reason is that $w_{p}$ is small in magnitude and it does not influence the nucleus radius.
IV.3 Main numerical results: Solution with curvature
@@
In the main calculation of this paper, we have incorporated the curvature correction to the surface tension, which is a contribution to the energy due to the principal curvature of a thin nuclear surface without finite-width corrections. As we have mentioned above, we use our self-consistent numerical results (as given in Table 3) on the leading (planar surface) contribution $\sigma_{s}(x)$ and the next-to-leading (curvature) correction $\sigma_{c}(x)$, for the surface tension of a thin surface. We represent these functions numerically with the help of spline interpolation – in this way we avoid choosing any fitting parameters such as in Eq. (9).
By following the scheme of solution of the variational equations explained above, we find the four basic variables, $\{n,x,r_{N},n_{no}\}$ (and the corresponding quantities for the bubble phases) as functions of mean (averaged across the WS cell) baryon density $n_{b}$ scaled by the saturation density in symmetric nuclear matter $n_{0}$ and display them in Figs. 1–4.
@@
Figures 1-4 show, correspondingly, the proton fraction $x$ inside the nucleus (or outside the bubble), the nucleus radius $r_{N}$ (or the scaled bubble radius, $r_{B}\times 10^{-1}$), the local baryon density $n$ inside the nucleus (or outside the bubble) and the local number density $n_{no}$ of neutrons outside of the nuclear cluster (or inside the bubble).
@@
Figure 3: Solution to the basic equations of equilibrium, Eqs. (4)-(7): The baryon number density inside the nucleus and outside the bubble as function of the mean baryon density $n_{b}$.
Figure 4: Solution to the basic equations of equilibrium, Eqs. (4)-(7): The baryon number density $n_{no}$ of pure neutron matter outside the nucleus and inside the bubble as function of the mean baryon density $n_{b}$.
IV.4 Proton workfunction
The variables shown in Figs. 1–4 represent a complete and unique solution to the basic equations, however, the appearance of pasta structures is not explicit. In order to reveal the appearance of pasta phases, as a first step, we compute the difference between the proton chemical potential inside and outside of the nuclear cluster (Eqs. (103), (159) and (104)),
@@ IV.4 Proton workfunction
which represents the workfunction of protons, which is a potential energy barrier created by the cluster, that prevents protons from flowing out of the nuclear cluster into the pure neutron matter.
@@
has been obtained assuming that the second phase is the pure neutron liquid. With this assumption, the surface energy is a function of the proton fraction only inside the nucleus (see Eqs. (73) and (73)). However, in the situation, not considered in this work, when the second phase is a dilute nuclear matter Vinas1998 ; douchin00 ; PethickEtal1995 ; KellerEtal2024 , the present liquid drop model should be extended in order to take into account nonzero proton fractions both inside and outside the nucleus.
V A simple model of superconductivity
In observable neutron stars, superconductivity of the pasta phases is important in the context of interaction between the inner crust and the stellar magnetic induction. A proper description should be capable to account for the smeared character of profiles of the nucleon densities and for other microscopic features of the pairing problem. Thus, the thermodynamic properties such as the critical magnetic field of the pasta phases, the proton-proton correlation length and the pairing energy, can be studied quantitatively. While this topic will be addressed in the future work, here, we introduce a simple model, which is expected to qualitatively grasp the basic physical properties without a need to develop a complicated theoretical framework.
In our model, the proton density profile is sharp (it is zero outside and finite and uniform inside the nuclear cluster). Similarly, the superconducting density is assumed to have a sharp profile. We evaluate the coherence length and the London penetration depth associated with matter inside the nuclear cluster using the results obtained in pairing calculations in uniform nuclear matter. The input data for our model is reported in Fig. 7 of LimHolt2021 for the proton pairing gap energy in beta equilibrium. Thus, energy gap has been obtained from the pairing calculations with the same underlying nuclear interaction, as has been used for the structure calculations.
The pairing energy contribution to the total energy is calculated from thermodynamic arguments. The excess (negative) energy associated with pairing is proportional to the gap energy and the number of paired particles at given temperature. This leads to the following expression for the pairing energy density in liquid drop model averaged over the Wigner-Seitz cell at low temperature:
$w_{p}=-\kappa unx\Delta_{p},\quad(\rm nucleus\;phases)$ (15)
$w_{p}^{\rm bub}=-\kappa(1-u^{\rm bub})nx\Delta_{p},\quad(\rm bubble\;phases)$ (16)
where $\kappa$ is a number that depends on microscopic properties and $\Delta_{p}$ is the value of the gap energy inside the nuclear cluster. The quantities appearing in Eqs. (15) and (16) are constant numbers and have no spatial dependence within the unit cell. Notice that the physical prediction for the pairing energy density that follows from the expression used by Pizzochero, Viverit and Broglia in Eq. (3) of Pizzochero1997 and later on by Pearson and Chamel in Eq. (14) of PearsonChamel2022 , is equivalent to the prediction following from our Eq. (15) with a certain choice for $\kappa$. This becomes evident as soon as Eq. (3) of Pizzochero1997 is spatially averaged over the Wigner-Seitz cell. It is easy to see that in this case
$\kappa=\frac{3}{8}\frac{\Delta_{p}}{E_{Fp}},$ (17)
where $E_{Fp}=\hbar^{2}(3\pi^{2}nx)^{2/3}/2{\tilde{m}}_{p}$ with ${\tilde{m}}_{p}$ given in Eq. (48).
VI Analysis of numerical results
VI.1 Pasta configurations with lowest internal energy and polymorphism
Here, we focus on two questions. (i) What is the pasta phase with the minimum internal energy? (ii) Which pasta phases have energy per baryon that is larger than the minimum by less than the thermal energy?
@@ VI.1 Pasta configurations with lowest internal energy and polymorphism
Inserting the solutions shown in Figs. 1–4 into the expressions for the energy density, Eqs. (70) and (130), we calculate the total internal energy density for each of the five pasta phases. Next, we determine the ground-state configuration at zero temperature, which is the one with the lowest internal energy density at fixed total baryon number and fixed temperature. Notice that for fixed $n_{b}$, as in our approach, the picture in terms of the energy per baryon is equivalent, up to a constant scaling factor, to the picture in terms of the energy density. As a next step, we find all the possible structures whose energy per baryon differs from the ground state by less than the thermal energy $k_{B}T$.
It is worth noting that to find the most thermodynamically favorable state of the system, one needs to minimize the thermodynamic potential $\Phi$. For this purpose, it is necessary to define a set of thermodynamic variables, which imposes restrictions on the form of $\Phi$. For instance, in this work the thermodynamic variables are the total number of baryons $N_{b}$ in the unit cell with volume $V_{c}$ and temperature $T$. In this case, $\Phi(N_{b},T)=E_{\rm tot}(N_{b})-k_{B}TS(N_{b},T)V_{c}$, where $E_{\rm tot}(N_{b})$ is the internal energy per unit cell and $S(N_{b},T)$ is the total entropy of the system per unit cell evaluated at given $N_{b}$ and $T$. Alternatively, an equivalent description can be given in terms of the total pressure $P$ and temperature $T$ and in this case $\Phi(P,T)=E_{\rm tot}(P)-k_{B}TS(P,T)V_{c}+PV_{c}$, where the internal energy and entropy are calculated for given $P$ and $T$. However, temperatures in the range of $10^{8}$$10^{9}$ K are very small compared to the typical nuclear energy scale of 1 MeV, thus, the relevant quantity is the internal energy at $T=0$. In this work we focus solely on the internal energy, while leaving the thermal effects for the future work.
@@
Our results are displayed in Fig. 6, where the minimum internal energy is shown by black solid line, while the dot markers show the solutions with energy per baryon that is larger than the minimum energy per baryon, by less than the thermal energy $k_{B}T$, for three different temperatures. The vertical blue lines show the critical densities $\eta_{p}^{*}$ and $\eta_{\rm uni}^{*}$ defined in Eq. (53) and in Fig. 5. The calculations for $n_{b}/n_{0}>0.6$ have been done on a sparser array of values since the uniform matter is the only possible solution in this case.
@@
Figure 6: (Color online) Thermodynamically allowed pasta configurations at three different temperatures $T$ taking on the values $10^{9}$ K (red), $5\times 10^{8}$ K (gold) or $10^{8}$ K (green). Black solid line shows the the ground-state characterized by the minimum internal energy per baryon calculated from Eqs. (71) and (131). The transitions between various pasta phases are seen as jumps of the black solid line, where the almost vertical pieces represent the absence of the data because of the finite step size of the numerical grid for $n_{b}$. The color markers indicate the excited configurations, which satisfy Eqs. (4)-(7), while their internal energy per baryon is larger than the minimum internal energy by less than $k_{B}T$. The dash-dotted line marks the critical value of $n_{b}$ at which uniform nuclear matter becomes unstable with respect to long wavelength hydrodynamic fluctuations (see Eq. (53) for details). The dashed line marks the critical value of $n_{b}$ at which protons start dripping out from the nuclei (see Fig. 5 for details).
Figure 6 suggests that even in the low-temperature regime at $T\sim 10^{8}$ K, the state of a real matter is likely a polymorphic structure, in which locally, nuclear clusters with various types of clusters (either of 1N, 2N, 3N, 2B and 3B or uniform) coexist. It is known from the material science that the process of crystallization might lead to simultaneous formation of various crystal structures with close value of the thermodynamic potential, ending up with a polymorphic rather than a pure crystal structure. A similar scenario for the crust was also anticipated in NewtonEtal2022 .
For instance, consider the case when $T=10^{8}$ K. As it follows from our calculations, for $n_{b}$ below $\sim 0.16n_{0}$, the system energetically favors only 3N phase. However, for $n_{b}$ above $\sim 0.16n_{0}$ and below $\sim 0.297n_{0}$, the system favors both 3N and 2N phases. For $n_{b}$ above $\sim 0.297n_{0}$ and below $\sim 0.345n_{0}$, or above $\sim 0.41n_{0}$ and below $\sim 0.515n_{0}$, the system favors 3N, 2N and 1N phases. For $n_{b}$ above $\sim 0.345n_{0}$ and below $\sim 0.409n_{0}$, the favored phases are 3N, 2N, 1N and 2B phases. Interestingly, for $n_{b}$ above $\sim 0.515n_{0}$ and below $n_{0}\eta_{\rm uni}^{*}$, the system favors 3N, 2N, 1N and the uniform phases.
VI.2 Characteristic physical lengths of protons
We start by noticing that Fig. 6 shows that lasagna is the ground state in a significant portion of the inner crust. In the weak pairing Bardeen-Cooper-Schrieffer approximation, the coherence length, which can be understood as the spatial size of the Cooper pair, reads